# Relatively big projective modules and their applications to direct sum decompositions

### v1

@@ 4. The case of right noetherian PI-rings
In this section we shall prove that countably generated, projective, right or left modules over a right noetherian PI-ring are always relatively big. The key result will be to show that they satisfy the ()(*)( ∗ ) condition introduced in §2. As a right noetherian PI-ring $R$ has nilpotent nilradical $N(R)$ and $R/N(R)$ is two-sided noetherian, the result will follow by combining Theorem 1.18 with the results in §3.
First we consider the case of commutative noetherian rings. P. Příhoda and G. Puninski proved in [PP] that commutative noetherian rings satisfy condition ()(*)( ∗ ). Here we give another proof of this result, which we find interesting as it is quite different from the one given in [PP].
@@ 4.1. Commutative noetherian rings
[…]
@@
Let $\Lambda$ be a ring. Let $(a_{k})_{k\geq 1}$ be a sequence of elements of $\Lambda$ satisfying $a_{k}a_{k+1}=a_{k+1}$ for any $k\geq 1$. Set $b_{k}=1-a_{k}$ for any $k\geq 1$. Then,
@@
    there is an descending sequence of principal right ideals
@@
    $\Lambda b_{k}\subseteq\mathrm{l.ann}(a_{k+1})\subseteq\Lambda b_{k+1}$
@@
    $I=\bigcup_{k\geq 1}\Lambda b_{k}$ is a left pure ideal of $R$ and it is also projective.
@@
    the sequence (4) is stationary if and only if so is the sequence (5). In this case, there exists $k_{0}$ such that $b_{k_{0}}$ and $a_{k_{0}}$ are idempotent, $I=\Lambda b_{k_{0}}$ and $\bigcap_{k\geq 1}a_{k}\Lambda=a_{k_{0}}\Lambda$.
Proof.  Statements $(i)$, $(ii)$ and $(iii)$ follow easily by direct computation.
@@
Statement $(iii)$ readily implies that (4) is stationary if and only if so is (5).
If $a_{k}\Lambda=a_{k+1}\Lambda$, then there exists $\lambda\in\Lambda$ such that $a_{k+1}\lambda=a_{k}$. Hence, $a_{k+1}\lambda a_{k+1}=a_{k+1}$ which implies that $a_{k}=a_{k+1}\lambda$ is idempotent.
@@
Proof.  Let $J=x_{1}R+\dots+x_{n}R$. The equality $JI=J$ implies that there exists $A\in M_{n}(I)$ such that $(x_{1},\dots,x_{n})A=(x_{1},\dots,x_{n})$ or, equivalently
@@
Multiplying the above identity on the right by the adjoint matrix of $\mathrm{Id}-A$, we deduce that $\det(A)$ annihilates $J$. To conclude the proof, note that there exists $a\in I$ such that $\det(A)=1-a$.
@@
([PP, Lemma 2.1]) Every commutative noetherian ring satisfies condition ()(*)( ∗ ). This is to say that if $R$ is a commutative noetherian ring and
@@
is a descending chain of ideals of $R$ satisfying $J_{k+1}J_{k}=J_{k+1}$ for each $k\geq 1$, then (6) is stationary.
In addition, in this context, $\bigcap_{k\geq 1}J_{k}=eR$ where $e$ is an idempotent of $R$.
Proof.  Let
$J_{1}\supseteq J_{2}\supseteq\cdots\supseteq J_{k}\supseteq\cdots$ (*)
be a descending chain of ideals of $R$ such that $J_{k+1}J_{k}=J_{k+1}$, for each $k\geq 1$.
Since $R$ is noetherian, the ideals $J_{k}$ are finitely generated for every $k\geq 1$. By Lemma 4.2, there exists a sequence of elements $(a_{k})_{k\geq 1}$ in $R$ such that
@@
is stationary and its union is generated by an idempotent of $R$. By Lemma 4.1, so is the descending sequence
@@
and its intersection is generated by an idempotent of $R$. The sequence (*4.1) is also stationary and the intersection of these ideals is generated by an idempotent of $R$.
In the following result, we recover a result of Bass that shows that, over commutative noetherian rings, the theory of countably generated projective modules reduces to the finitely generated ones.
@@
[bass] Let $R$ be a commutative noetherian ring. Let $P$ be a countably generated projective module that is not finitely generated. Then there exists an idempotent $e\in R$ with $P\cong(1-e)P\oplus\left(eR\right)^{(\omega)}$ and such that $(1-e)P$ is finitely generated.
Proof.  Let $P$ be a countably generated projective module over $R$. By Proposition 4.3 and Theorem 1.18, there exists an idempotent $e\in R$ such that $P/eP$ is finitely generated, $I=eR$ is an ideal minimal with respect to that property, and $P$ is $I$-big. Therefore $P\cong P\oplus(eR)^{(\omega)}$.
@@
where the last isomorphism follows by the Eilenberg swindle. To conclude the proof of the statement, observe that $(1-e)P\cong P/eP$.
@@
Not much is known on the structure of infinitely generated projective modules over general commutative domains. To our knowledge, the basic questions on the topic posed in [fuchssalce, p.  246] are still widely open and it seems that to answer them one needs to go beyond the theory of relatively big projectives.
In general, over a commutative domain $R$, the only trace ideals are $\{0\}$ and $R$, so $B(R)$ contains only the isomorphism class of $R^{(\omega)}$. Therefore, $V(R)\sqcup B(R)=V(R)\sqcup\{[R^{(\omega)}]\}$. We do not know any reasonable characterization of domains $R$ such that $V^{*}(R)=V(R)\sqcup B(R)$, neither examples where this fails.
@@ 4.2. One-sided noetherian PI-rings
By the results in [P3] or in [wiegand] that the structure of countably projective modules over finitely generated algebras over a commutative noetherian ring is more involved than the one described in Corollary 4.4. The results on these papers were depending on the fact that if $\Lambda$ is a finitely generated algebra over a commutative noetherian ring of Krull dimension $1$ then $V^{*}(\Lambda)=V(\Lambda)\sqcup B(\Lambda)$, that is, all countably generated projective, right or left, $\Lambda$-modules are relatively big. In this subsection we show this is true for countably generated projective modules over any right noetherian, PI-ring $\Lambda$.
@@
Proof.  Using the hypothesis we deduce that,
@@
Assume $JI=J$,then $JI^{n}=J$. Since $I^{n}\subseteq J$, we deduce that $J\subseteq J^{2}$. This shows that $J=J^{2}$.
@@
    Let $N(\Lambda)$ denote the prime radical of $\Lambda$ (that is, the intersection of all prime ideals of $\Lambda$). If (7) or (8) are stationary in $\Lambda/N(\Lambda)$ then they are also stationary over $\Lambda$.
@@
We shall prove in Theorem 4.8 that, in addition, right noetherian PI-rings satisfy condition ()(*)( ∗ ) and its left analogue that we call ()(**)( ∗ ∗ ). The proof is patterned on the one by Robson and Small.
First we give some basic definitions needed to follow the proof.
@@
A ring $\Lambda$ is said to be a polynomial identity ring, a PI-ring for short, if there exists $n$ such that $\Lambda$ satisfies a monic polynomial of $\mathbf{Z}\langle x_{1},\dots,x_{n}\rangle$. That is, there exists a polynomial in non-commuting variables $x_{1},\dots,x_{n}$ with coefficients in $\mathbf{Z}$, $f(x_{1},\dots,x_{n})$ say, such that:
@@
    the coefficient of one of the monomials of highest degree of $f$ is one.
@@
where $\varepsilon(\sigma)$ denotes the sign of the permutation $\sigma$. For example, if $R$ is a non-zero commutative ring, a theorem of Amitsur and Levitzki shows that $M_{n}(R)$ satisfies the standard identity $s_{m}$ for any $m\geq 2n$ and does not satisfy $s_{2n-1}$, cf. [MR, §13.3.2].
@@
It is possible to define the PI-degree of a central simple algebra $\Lambda$ (that is, a simple algebra that is finite dimensional over its center) as the minimum $n$ such that $\Lambda$ satisfies $s_{2n}$ and not $s_{2n-1}$. As prime PI-rings localized at the non-zero elements of the center are central simple algebras, this definition of the PI-degree can be extended to prime PI-rings. For details, check [MR, §13.6.7].
@@
Let $\Lambda$ be a right noetherian PI-ring. Then $\Lambda$ satisfies condition ()(*)( ∗ ) and its symmetric ()(**)( ∗ ∗ ).
Proof.  Assume the statement is not true. By noetherian induction, we may assume that $\Lambda$ is a right noetherian PI-ring that does not satisfy condition ()(*)( ∗ ) but so does any proper quotient ring of $\Lambda$.
By [RS, Theorem 4], $\Lambda$ has only a finite number of idempotent ideals. By Proposition 4.6 this implies that $\Lambda$ is a prime, right noetherian, PI-ring. Therefore, $\Lambda$ is (right and left) noetherian by a theorem of G. Cauchon [MR, 13.6.15 Theorem]. Now we adapt the steps of the proof of [RS, Theorem 4] to get a contradiction.
@@
be a sequence of ideals of $\Lambda$ satisfying that $I_{k+1}I_{k}=I_{k+1}$ for any $k\geq 1$. Let $P$ be a prime ideal containing $I_{k}$ for some $k$ and assume that $\mathrm{deg}\,\Lambda/P=\mathrm{deg}\,\Lambda$. By [small, Theorem], the set of elements which are regular modulo $P$ consists of regular elements of $\Lambda$, forms an Ore set $S$ and the localization $\Lambda S^{-1}$ is a local ring with maximal ideal $P\Lambda S^{-1}$. Since $I_{k}$ is contained in $P$, so is $I_{k+\ell}$ for any $\ell\geq 1$. But in the local ring $\Lambda S^{-1}$, $P\Lambda S^{-1}I_{k+\ell}\Lambda S^{-1}=I_{k+\ell}\Lambda S^{-1}$ for any $\ell\geq 1$. As the ring is noetherian, Nakayama’s Lemma implies that $I_{k+\ell}\Lambda S^{-1}=\{0\}$ for any $\ell\geq 1$. Therefore, $I_{k+\ell}=\{0\}$ for any $\ell\geq 1$. This contradicts our hypothesis. Therefore, $\mathrm{deg}\,\Lambda/P<\mathrm{deg}\,\Lambda$ for any prime ideal $P$ containing $I_{k}$ for some $k$.
Let $n=\mathrm{deg}\,\Lambda$. Let $E(\Lambda)$ be the nonzero ideal generated by the evaluations in $\Lambda$ of $s_{2n-2}$ (cf. [MR, 13.6.7 Corollary]). We have just proved that for any fixed $k$, and for any prime ideal $P$ containing $I_{k}$, the ideal $E(\Lambda)$ is contained in $P$. Therefore, $E(\Lambda)$ is contained in the prime radical of $I_{k}$. Therefore, for any $k$ there exists $m_{k}$ such that $E(\Lambda)^{m_{k}}\subseteq I_{k}$ (recall that the prime radical of $\Lambda/I_{k}$ is nilpotent by [GW, Theorem 2.11]).
By hypothesis, $\Lambda/E(\Lambda)$ satisfies condition ()(*)( ∗ ). This implies that there exists $k_{0}$ such that $I_{k_{0}+\ell}+E(\Lambda)=I_{k_{0}}+E(\Lambda)$ for any $\ell\geq 0$. By Lemma 4.5, we deduce that $I_{k_{0}+\ell+1}$ is an idempotent ideal for any $\ell\geq 0$. As a consequence, the sequence (9) contains infinitely many different idempotent ideals, which gives the desired contradiction with the statement of [RS, Theorem 4]. This concludes the proof of the statement.
@@
    Every idempotent ideal of $\Lambda$ is the trace of a countably generated projective right $\Lambda$-module and of a countably generated projective left $\Lambda$-module.
@@
Proof.  Statement $(i)$ follows from the fact that $\Lambda/N(\Lambda)$, where $N(\Lambda)$ denotes the prime radical of $\Lambda$, is two-sided noetherian [MR, 13.6.15], and that $N(\Lambda)$ is nilpotent. Therefore, if $I$ is an idempotent ideal of $\Lambda$, then $(I+N(\Lambda))/N(\Lambda)$ is an idempotent ideal of $\Lambda/N(\Lambda)$, so it is the trace of a countably generated projective right $\Lambda/N(\Lambda)$-module and the trace of a countably generated projective left $\Lambda/N(\Lambda)$-module (cf. [whitehead]). By Lemma 3.1, $I$ is also the trace ideal of a countably generated projective module on both sides.
By Theorem 4.8, $\Lambda$ and $\Lambda/N$ satisfy the ()(*)( ∗ )-condition and the symmetric ()(**)( ∗ ∗ )-condition. By Theorem 2.13, $V^{*}(\Lambda/N)=V(\Lambda/N)\sqcup B(\Lambda/N)$ and from Corollary 3.3 we deduce that $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}^{*}(\Lambda)=V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$.
Since $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$, the rest of the statement follows from Corollary 2.12.
It is an interesting open question to determine whether there exists a right noetherian ring $\Lambda$ with an idempotent ideal that is the trace of a (countably generated) projective, left $\Lambda$-module but not the trace of a projective right $\Lambda$-module. Corollary 4.9 shows that if such an example exists, it cannot be a PI-ring.
The following corollary was our initial motivation to develop these results. As it shows that the basics of the machinery developed in [wiegand] for the endomorphism ring of finitely generated modules over commutative noetherian rings of Krull dimension $1$, is also available in arbitrary Krull dimension.
@@
Let $R$ be a commutative noetherian ring $R$, and let $\Lambda$ be a module-finite $R$-algebra. Then $\Lambda$ satisfies condition ()(*)( ∗ ) and its symmetric condition ()(**)( ∗ ∗ ). Moreover, it has only a finite number of idempotent ideals, and projective modules satisfy the conclusions of Corollary 4.9.
Proof.  If $\Lambda$ is generated by $n$-elements as $R$-module, then $\Lambda$ is a sub-quotient of $M_{n}(R)$, so it is PI and noetherian (cf. [MR, 13.4.9 Corollary]). Then, Theorem 4.8 implies that $\Lambda$ satisfies condition ()(*)( ∗ ).

### v2

@@ 4. The case of right noetherian PI-rings
In this section, we shall prove that countably generated, projective, right or left modules over a right noetherian PI-ring are always relatively big. The key result will be to show that these rings satisfy the ()(*)-condition introduced in §2. As a right noetherian PI-ring $R$ has nilpotent nilradical $N(R)$, and $R/N(R)$ is two-sided noetherian, the result will follow by combining Theorem 1.23 with the results in §3.
First, we consider the case of commutative noetherian rings. P. Příhoda and G. Puninski proved in [PP] that commutative noetherian rings satisfy condition ()(*). Here, we provide another proof of this result, which we find interesting as it is quite different from the one given in [PP].
@@ 4.1. Commutative noetherian rings
If $I$ is a countably generated left ideal of a ring $\Lambda$, $I$ is the union of an ascending sequence of the type (5) in Lemma 4.1 if and only if $I$ is a pure left ideal [lazard, Lemme 2].
Over a commutative ring, an ideal is the trace of a projective module if and only if it is pure (see Remark 1.7). Since the trace ideal of a countably generated projective module is countably generated, an ideal is the trace of a countably generated projective module if and only if it is the union of an ascending sequence of the type (5) in Lemma 4.1.
@@
Let $\Lambda$ be a ring. Let $(a_{k})_{k\geq 1}$ be a sequence of elements of $\Lambda$ satisfying $a_{k}a_{k+1}=a_{k+1}$ for any $k\geq 1$. Set $b_{k}=1-a_{k}$ for any $k\geq 1$. Then:
@@
    There is a descending sequence of principal right ideals
@@
    $\Lambda b_{k}\subseteq\mathrm{l.ann}(a_{k+1})\subseteq\Lambda b_{k+1},$
@@
    $I=\bigcup_{k\geq 1}\Lambda b_{k}$ is a left pure ideal of $R$, and it is also projective.
@@
    The sequence (4) is stationary if and only if the sequence (5) is also stationary. In this case, there exists $k_{0}$ such that $b_{k_{0}}$ and $a_{k_{0}}$ are idempotent, $I=\Lambda b_{k_{0}}$ and $\bigcap_{k\geq 1}a_{k}\Lambda=a_{k_{0}}\Lambda$.
Proof. Statements $(i)$, $(ii)$, and $(iii)$ follow easily from direct computation.
@@
Statement $(iii)$ readily implies that (4) is stationary if and only if (5) is also stationary.
If $a_{k}\Lambda=a_{k+1}\Lambda$, then there exists $\lambda\in\Lambda$ such that $a_{k+1}\lambda=a_{k}$. Hence, $a_{k+1}\lambda a_{k+1}=a_{k+1}$, which implies that $a_{k}=a_{k+1}\lambda$ is idempotent.
@@
Proof. Let $J=x_{1}R+\dots+x_{n}R$. The equality $JI=J$ implies that there exists $A\in M_{n}(I)$ such that $(x_{1},\dots,x_{n})A=(x_{1},\dots,x_{n})$, or equivalently
@@
Multiplying the above identity on the right by the adjoint matrix of $\mathrm{Id}-A$, we deduce that $\mathrm{Id}-\det(A)$ annihilates $J$. To conclude the proof, note that there exists $a\in I$ such that $\det(\mathrm{Id}-A)=1-a$.
@@
([PP, Lemma 2.1]) Every commutative noetherian ring satisfies condition ()(*). This is to say that if $R$ is a commutative noetherian ring and
@@
is a descending chain of ideals of $R$ satisfying $J_{k+1}J_{k}=J_{k+1}$ for each $k\geq 1$, then (6) is stationary. In addition, $\bigcap_{k\geq 1}J_{k}=eR$ where $e$ is an idempotent of $R$.
Proof. Since $R$ is noetherian, the ideals $J_{k}$ are finitely generated for every $k\geq 1$. By Lemma 4.2, there exists a sequence of elements $(a_{k})_{k\geq 1}$ in $R$ such that
@@
is stationary and its union is generated by an idempotent of $R$. By Lemma 4.1, the descending sequence
@@
is also stationary, and its intersection is generated by an idempotent of $R$. The sequence (6) is also stationary, and the intersection of these ideals is generated by an idempotent of $R$.
In the following result, we recover a result of Bass that shows that, over commutative noetherian rings, the theory of countably generated projective modules reduces to that of finitely generated ones.
@@
[bass] Let $R$ be a commutative noetherian ring. Let $P$ be a countably generated projective module that is not finitely generated. Then there exists an idempotent $e\in R$ with $P\cong(1-e)P\oplus\left(eR\right)^{(\omega)}$ such that (1e)P(1-e)P is finitely generated.
Proof. Let $P$ be a countably generated projective module over $R$. By Proposition 4.3, Lemma 1.20, and Theorem 1.23, there exists an idempotent $e\in R$ such that $P/eP$ is finitely generated, $I=eR$ is an ideal minimal with respect to that property, and $P$ is $I$-big. Therefore $P\cong P\oplus(eR)^{(\omega)}$.
@@
where the last isomorphism follows from the Eilenberg swindle. To conclude the proof of the statement, observe that $(1-e)P\cong P/eP$.
@@
Not much is known about the structure of infinitely generated projective modules over general commutative domains. To our knowledge, the basic questions on the topic posed in [fuchssalce, p.  246] are still widely open, and it seems that to answer them, one needs to go beyond the theory of relatively big projectives.
In general, over a commutative domain $R$, the only trace ideals are $\{0\}$ and $R$, so $B(R)$ contains only the isomorphism class of $R^{(\omega)}$. Therefore, $V(R)\sqcup B(R)=V(R)\sqcup\{[R^{(\omega)}]\}$. We do not know of any reasonable characterization of domains $R$ such that $V^{*}(R)=V(R)\sqcup B(R)$, nor are there any examples where this fails.
@@ 4.2. One-sided noetherian PI-rings
In this subsection, we show that countably generated projective right or left modules over any right noetherian, PI-ring $\Lambda$ are relatively big.
The big difference between a noetherian PI-ring $\Lambda$ and the commutative case described in Corollary 4.4 is that the idempotent ideals are not just the ideals generated by idempotent elements of $\Lambda$. By a result of Small and Robson [RS], the set of idempotent ideals of $\Lambda$ (recall that it coincides with $\mathcal{T}(\Lambda)$) is finite. But the results from [P3] and [wiegand] provide plenty of examples showing that, even in the case where $\Lambda$ is a finitely generated algebra over a commutative noetherian ring, $\mathcal{T}(\Lambda)$ frequently contains elements that are not traces of finitely generated projective modules.
@@
Proof. Using the hypothesis, we deduce that,
@@
If $JI=J$, then $JI^{n}=J$. Since $I^{n}\subseteq J$, we deduce that $J\subseteq J^{2}$. This shows that $J=J^{2}$.
@@
    Let $N(\Lambda)$ denote the prime radical of $\Lambda$ (that is, the intersection of all prime ideals of $\Lambda$). If (7) or (8) are stationary in $\Lambda/N(\Lambda)$, then they are also stationary over $\Lambda$.
@@
We shall prove in Theorem 4.8 that, in addition, right noetherian PI-rings satisfy condition ()(*) and its left analog, which we call ()(**). The proof is patterned after that of Robson and Small.
First, we provide some basic definitions needed to follow the proof.
@@
A ring $\Lambda$ is said to be a polynomial identity ring, a PI-ring for short, if there exists $n$ such that $\Lambda$ satisfies a monic polynomial of $\mathbf{Z}\langle x_{1},\dots,x_{n}\rangle$. That is, there exists a polynomial in non-commuting variables $x_{1},\dots,x_{n}$ with coefficients in $\mathbf{Z}$, say $f(x_{1},\dots,x_{n})$, such that:
@@
    the coefficient of one of the monomials of the highest degree of $f$ is one.
@@
where $\varepsilon(\sigma)$ denotes the sign of the permutation $\sigma$. For example, if $R$ is a non-zero commutative ring, a theorem of Amitsur and Levitzki shows that $M_{n}(R)$ satisfies the standard identity $s_{m}$ for any $m\geq 2n$ and does not satisfy $s_{2n-1}$, cf. [MR, §13.3.2 and §13.3.2]. Moreover, if a PI-ring satisfies $s_{m}$, then it satisfies $s_{t}$ for any $t\geq m$, cf. [MR, §13.3.3].
@@
It is possible to define the PI-degree of a central simple algebra $\Lambda$ (that is, a simple algebra that is finite dimensional over its center) as the minimum $n$ such that $\Lambda$ satisfies $s_{2n}$ and not $s_{2n-1}$. In this case, we write $\mathrm{PI\mbox{-}deg}(\Lambda)=n$.
Prime PI-rings $\Lambda$ localized at the non-zero elements of the center are central simple algebras. Then, by definition, the PI-degree of $\Lambda$ is the PI-degree of this localization. For details, check [MR, §13.6.7].
@@
Let $\Lambda$ be a right noetherian PI-ring. Then $\Lambda$ satisfies condition ()(*) and its symmetric ()(**).
Proof. Assume the statement is not true. By noetherian induction, we may assume that $\Lambda$ is a right noetherian PI-ring that does not satisfy condition ()(*), but so does any proper quotient ring of $\Lambda$.
By [RS, Theorem 4], $\Lambda$ has only a finite number of idempotent ideals. By Proposition 4.6, this implies that $\Lambda$ is a prime, right noetherian, PI-ring. Therefore, $\Lambda$ is (right and left) noetherian by a theorem of G. Cauchon [MR, 13.6.15 Theorem]. Now we adapt the steps of the proof of [RS, Theorem 4] to obtain a contradiction.
@@
be a sequence of ideals of $\Lambda$ satisfying that $I_{k+1}I_{k}=I_{k+1}$ for any $k\geq 1$. Let $P$ be a prime ideal containing $I_{k}$ for some $k$ and assume that $\mathrm{PI\mbox{-}deg}\,(\Lambda/P)=\mathrm{PI\mbox{-}deg}\,(\Lambda)$. By [small, Theorem], the set of elements that are regular modulo $P$ consists of regular elements of $\Lambda$, forms an Ore set $S$, and the localization $\Lambda S^{-1}$ is a local ring with maximal ideal $P\Lambda S^{-1}$. Since $I_{k}$ is contained in $P$, so is $I_{k+\ell}$ for any $\ell\geq 1$. But in the local ring $\Lambda S^{-1}$, $P\Lambda S^{-1}I_{k+\ell}\Lambda S^{-1}=I_{k+\ell}\Lambda S^{-1}$ for any $\ell\geq 1$. As the ring is noetherian, Nakayama’s lemma implies that $I_{k+\ell}\Lambda S^{-1}=\{0\}$ for any $\ell\geq 1$. Therefore, $I_{k+\ell}=\{0\}$ for any $\ell\geq 1$. This contradicts our hypothesis. Therefore, $\mathrm{PI\mbox{-}deg}\,(\Lambda/P)<\mathrm{PI\mbox{-}deg}\,(\Lambda)$ for any prime ideal $P$ containing $I_{k}$ for some $k$.
Let $n=\mathrm{PI\mbox{-}deg}\,(\Lambda)$. Let $E(\Lambda)$ be the non-zero ideal generated by the evaluations in $\Lambda$ of $s_{2n-2}$ (cf. [MR, 13.6.7 Corollary]). For any fixed $k$, and for any prime ideal $P$ containing $I_{k}$, the ideal $E(\Lambda)$ is contained in $P$ because $\mathrm{PI\mbox{-}deg}\,(\Lambda/P)<\mathrm{PI\mbox{-}deg}\,(\Lambda)$. Thus, $\Lambda/P$ must satisfy $s_{2n-2}$. Therefore, $E(\Lambda)$ is contained in the prime radical of $I_{k}$. Then, for any $k$, there exists $m_{k}$ such that $E(\Lambda)^{m_{k}}\subseteq I_{k}$ (recall that the prime radical of $\Lambda/I_{k}$ is nilpotent by [GW, Theorem 3.11]).
By hypothesis, $\Lambda/E(\Lambda)$ satisfies condition ()(*). This implies that there exists $k_{0}$ such that $I_{k_{0}+\ell}+E(\Lambda)=I_{k_{0}}+E(\Lambda)$ for any $\ell\geq 0$. By Lemma 4.5, we deduce that $I_{k_{0}+\ell+1}$ is an idempotent ideal for any $\ell\geq 0$. As a consequence, the sequence (9) contains infinitely many different idempotent ideals, which gives the desired contradiction with the statement of [RS, Theorem 4]. This concludes the proof of the statement.
@@
    Every idempotent ideal of $\Lambda$ is the trace of a countably generated projective right $\Lambda$-module and the trace of a countably generated projective left $\Lambda$-module.
@@
Proof. Statement $(i)$ follows from the fact that $\Lambda/N(\Lambda)$, where $N(\Lambda)$ denotes the prime radical of $\Lambda$, is two-sided noetherian [MR, 13.6.15], and that $N(\Lambda)$ is nilpotent. Therefore, if $I$ is an idempotent ideal of $\Lambda$, then $(I+N(\Lambda))/N(\Lambda)$ is an idempotent ideal of $\Lambda/N(\Lambda)$, so it is the trace of a countably generated projective right $\Lambda/N(\Lambda)$-module and the trace of a countably generated projective left $\Lambda/N(\Lambda)$-module (cf. [whitehead]). By Lemma 3.1, $I\in\mathcal{T}_{r}(\Lambda)$, and $I\in\mathcal{T}_{\ell}(\Lambda)$.
By Theorem 4.8, $\Lambda$ and $\Lambda/N(\Lambda)$ satisfy the ()(*)-condition and the symmetric ()(**)-condition. By Theorem 2.10, $V^{*}(\Lambda/N(\Lambda))=V(\Lambda/N(\Lambda))\sqcup B(\Lambda/N(\Lambda))$, and from Corollary 3.3, we deduce that $V_{r}^{*}(\Lambda)=V_{r}(\Lambda)\sqcup B_{r}(\Lambda)$ and $V_{\ell}^{*}(\Lambda)=V_{\ell}(\Lambda)\sqcup B_{\ell}(\Lambda)$.
Since, by $(i)$, $\mathcal{T}_{r}(\Lambda)=\mathcal{T}_{\ell}(\Lambda)$, the rest of the statement follows from Corollary 2.9.
It is an interesting open question to determine whether there exists a right noetherian ring $\Lambda$ with an idempotent ideal that is the trace of a (countably generated) projective left $\Lambda$-module but not the trace of a projective right $\Lambda$-module. Corollary 4.9 shows that if such an example exists, it cannot be a PI-ring.
The following corollary was our initial motivation for developing these results. As it shows, the basics of the machinery developed in [wiegand] for the endomorphism ring of finitely generated modules over commutative noetherian rings of Krull dimension $1$ are also available in arbitrary Krull dimension. See also §4.3 for the transfer of information between direct sum decompositions of a finitely generated module and a projective module over its endomorphism ring.
@@
Let $R$ be a commutative noetherian ring $R$, and let $\Lambda$ be a module-finite $R$-algebra. Then $\Lambda$ satisfies condition ()(*) and its symmetric condition ()(**). Moreover, it has only a finite number of idempotent ideals, and projective modules satisfy the conclusions of Corollary 4.9.
Proof. If $\Lambda$ is generated by $n$-elements as an $R$-module, then $\Lambda$ is a sub-quotient of $M_{n}(R)$, so it is PI and noetherian (cf. [MR, 13.4.9 Corollary]). Then, Theorem 4.8 implies that $\Lambda$ satisfies the ()(*) condition and the symmetric ()(**) condition.